OpenStudy (anonymous):

Rob was negotiating a new contract with is team. He wanted $700,000 for the year and an additional $1800 for each game. His team was offering him $600,000 for the year, and 5000 for every game started. How many starts would he need to make in order to earn more with the teams offer? PLEASE HELP ME!

OpenStudy (anonymous):

Let x = # of games 1st contract: 700000 + 1800x 2nd contract: 600000 + 5000x Set the two equal to each other to find out the number of games where the contracts are equal. 700000 + 1800x = 600000 + 5000x 100000 = 3200x x = 31.25 games The answer needs to be rounded up in a whole number of games, so 32.